# A systematic comparison of measures for k-anonymity in networks

### v1

@@ 4.3.4 Hybrid
The previously mentioned measures can also be combined, which is done in [67]. Here nodes are equivalent if they have isomorphic 1-neighborhoods and if the degree distribution over the neighboring nodes is equal. This is referred to as the $1^{*}$-neighborhood attack. Hence, this combines the measures of 1-neighborhood isomorphism [11, 45, 65, 12, 73], and VRQ [38, 27]. A parameterized version would combine the $d$-neighborhood with VRQ$(d)$.
4.3.5 Anonymity-cascade
@@
Recent work introduces anonymity-cascade [12]. This algorithm models the scenario that an attacker reuses the nodes that are uniquely identified to identify more nodes in the network. This is similar to the attacker scenarios that include a propagation step [48, 49] as discussed in Section 2.2. In the work of [12], the algorithm starts with all nodes that have a unique $1$-neighborhood structure. Then, it continues in a cascading fashion by identifying neighboring nodes with a unique structure as unique. This way the reach of the measure is extended beyond the direct neighborhood.
4.3.6 Other structural knowledge
@@
For completeness of the summary, we end with approaches that take into account different structural information, uncertainty and other graph properties. Several works introduced measures that account for other structural information than discussed thusfar. We summarize four such methods being edge facts, hub fingerprints, a method based on centrality of nodes, and one based on twin nodes.
@@
First, we look at approaches introduced in the work of Hay et al. [38, 27]. Besides vrq, it introduces two different scenarios for which the first scenario is based on knowledge about edge facts. This approach measures anonymity by counting how many candidates there are given that an attacker knows a number of edges surrounding the node. However, computing all possible subgraphs given a number of edges is very computationally expensive and would not scale to large graphs. At the same time, if the edges are limited to the $d$-neighborhood of the graph, it can be at most as strict as the distance-parameterized structural neighborhood measures mentioned at the end of Section 4.3.2.
@@
The second approach is the hub fingerprints approach. In this approach, it is assumed that the attacker knows the distance for each node to so-called hub nodes. These are nodes that have a central position in the network and often have a high degree. However, experiments showed that this knowledge helps to identify only few nodes. As this is not an effective attacker scenario, we chose to not include this measure in our comparison.
@@
Third, the work of [40] introduces an attacker scenario where the attacker has additional structural information besides the neighborhood of a node. Namely, this work assumes the attacker can obtain information about the centrality of nodes and explicitly aims to protect against this scenario.
@@
Fourth, as part of a different approach, [12] introduces the notion of twin-uniqueness. Under the assumption that the graph is shared, the attacker can identify so called twin nodes, which are sets of nodes that are connected to the same nodes. When an attacker finds that all candidates for an entity of interest are twin, they can learn all structural information there is to know about the node representing the entity, including the connections. The notion of twin-uniqueness accounts for this phenomenon when computing anonymity and adds the constraint that for a node to be anonymous at least one equivalent node should not be a twin.
4.3.7 Uncertainty
The work in [75] builds upon the notion of $k$-degree anonymity [26, 68, 69, 70, 64, 43, 44] and assumes there might be uncertainty in the knowledge of the attacker. This is accounted for by deploying a so-called binning approach such that nodes with a similar degree, which are in the same bin, are also equivalent. Since this approach accounts for uncertainty, it can merge equivalence classes that are distinct according to degree, hence it is as most as strict as this measure.
4.3.8 Network model extensions
Rather than focusing on undirected networks, of course there has been work that accounts for more network properties. Various works extend the notion of $k$-anonymity to include node labels [51, 72, 73, 76], edge labels [77] or edge weights [78]. Another work focuses on weighted dynamic networks with node labels [69]. Since all of these properties add more information, extending an anonymity measure to account for both network structure and such additional properties results in a situation in which the number of unique nodes found will always be higher, or in the best case equal. Hence, accounting for additional properties overall decreases the anonymity.
@@ 5 Theoretical comparison of k-anonymity network measures
In the previous section, we summarized the anonymity measures introduced in literature. These measures differ in how complete the structural information is they take into account, and their reach, i.e., how far from the target node this structural information is taken into account. In this section we aim to understand the most prominent measures for anonymity introduced in the literature We do so by summarizing their reach and the structural information they take into account allowing us to order them based on strictness.
@@
In our comparison, we account for the measures included in Table 3. All measures included, except for degree, are parameterized by parameter $d$. This value determines up to what distance the structural information is taken into account from the target node. For these parameterized variants, the equivalence classes are computed recursively. Classes are split into new classes as the distance increases. Below, we briefly summarize each of the measures used. An equation and the abbreviation used throughout this manuscript for each measure can be found in Table 3.
@@
  • count: the number of nodes and edges in the $d$-neighborhood of the considered node.
  • degdist: the degree distribution of the $d$-neighborhood of the considered node. Note that this is a multiset.
@@
  • vrq: the degree of all nodes at distance at most $d$ from the considered node. Note that this is a multiset.
  • hybrid: set containing the results of $d$-$k$-anonymity and vrq, i.e., the exact structure of the $d$-neighborhood of the considered node summarized by its canonical labeling and the degree of all nodes at distance at most $d$ from the considered node.
For some measures nodes can only be equivalent if they are also equivalent for another measure. This is for example the case for count and degree. If the nodes are equivalent for count, they also need to be equivalent for degree. It follows that if the nodes are not equivalent for degree, they can not be equivalent for count. We define this notion as strictness in Definition 5. Overall, when a measure is more strict than another measure, more information should be taken into account compared to the less strict measure. As a result, more nodes may be distinguishable compared to when a more lenient measure is used. Hence, the use of the stricter measure results in more equivalence classes that are smaller in size, and thus lower overall network anonymity.
Definition 5 (Strictness)
Given two measures $M_{1}$ and $M_{2}$, we say that $M_{2}$ is more strict than $M_{1}$, denoted $M_{1}\leq M_{2}$, if the following holds for each pair of nodes $v,w$:
  • $v\cong_{M_{2}}w\rightarrow v\cong_{M_{1}}w$
We can use this notion to order the measures for $k$-anonymity. First, we look at the neighborhood based measures being degree, count, degdist and $d$-$k$-anonymity. As illustrated in the second and fourth column of Table 2, these measures all have the same reach, namely the $d$-neighborhood of the node. However, the measures do differ in the extent to which they capture particular structural properties. Theorem 1 defines how these measures relate to each other based on strictness. We prove this ordering holds by showing that the stricter measures capture properties of the less strict measures. The ordering resulting from Theorem 1 is visualized in Figure 1.
Figure 1: Ordering of anonymity measures based on strictness. An arrow between two measures ($A\rightarrow B$) implies measure $A$ is more strict than measure $B$ ($B\leq A$).
Theorem 1
For $d\geq 1$, the following holds:
  • degree $\leq$ count$(d)$leq $(d)$leq $d$-$k$-anonymity $\leq$ hybrid$(d)$
  • vrq$(d)$leq hybrid$(d)$
For $d\geq 2$:
  • vrq$(d-1)$leq $d$-$k$-anonymity.
  • vrq$(d-1)$leq vrq$(d)$.
Proof 1
Given a graph $G=(V,E)$ and nodes $v,w$:
  • degree $\leq$ count. Since the degree of the target node equals the number of nodes in the $1$-neighborhood minus one, the degree can be derived from count at $d=1$. As a result, two nodes equivalent according to count are equivalent for degree, ensuring the requirement: $v\cong_{\textsc{count}}w\rightarrow v\cong_{\textsc{degree}}w$.
Theorem 2
[…]
Proof 2
We prove this theorem by use of the counterexamples in Figure 2. The figure shows one case for which count is able to distinguish between two nodes and vrq is not, an example showing the converse where vrq is able to distinguish between the nodes and count is not and an example where vrq is able to distinguish between two nodes and count$(d)$ and $d+1$ and is not. Hence, we can not conclude that one measure is more strict than the other.
@@ 6 Empirical comparison of k-anonymity measures
In the upcoming sections, we empirically compare the anonymity measures in Table 3 on a wide range of network datasets obtained from various online repositories listed in Table 4. First, we summarize the experimental setup. Then we compare uniqueness on the datasets and thereafter the runtime required to compute anonymity for each measure. Lastly, we compare uniqueness of anonymity-cascade.
6.1 Experimental setup for empirical comparison
First, we investigate anonymity in terms of uniqueness, i.e., the fraction of nodes that are $1$-anonymous using various measures. We look at results for both $d=1$ and $d=2$, as previous work found a relatively large increase in uniqueness for $d$-$k$-anonymity with $d=2$ and relatively small for $d>2$ [12]. To compute uniqueness, we use the same computational approach as described in [65]. Second, for the runtime of the different measures, we report results averaged over 5 runs with a time limit of 3 hours. Runs longer than 3 hours are not reported. Code used can be found in the following repository [^1] https://github.com/RacheldeJong/ANONET.
Network Nodes Edges
Average
@@ 6.1 Experimental setup for empirical comparison
Table 4: Real-world network datasets used in the experiments. For each network, we list the number of nodes, edges, average degree, median degree, largest degree, clustering coefficient, assortativity, diameter, average distance and alpha value.
Third, we measure the uniqueness when using a more generic version of anonymity-cascade [12]. For this generic version, we distinguish between the initial measure and cascading measure. We report results on all 36 combinations that can be made using the six measures in Table 3.
@@ 6.2 Comparing uniqueness
From Theorem 1, we know that some considered measures are stricter than others. Hence, the uniqueness is expected to be higher for stricter measures and lower for less strict measures. Figure 3 shows how the uniqueness varies in empirical networks by reporting this value for each measure at both $d=1$ and $d=2$. In Appendix A, we report for each measure the difference with the previous measure in the legend of Figure 3. In an attempt to understand performance differences, we compare each difference with various topological network properties, of which the results can be found in Table 5. This table shows for each combination of difference and network property the Pearson correlation and $p$-value. Plots of the values for all significantly correlated combinations can be found in Appendix B. We use these findings in the discussion of the results.
@@
Table 5: Correlation between network properties (leftmost column) and the difference observed between the use of two measures (second up to last column). For each combination of network property and outcome both the Pearson correlation and $p$-value are reported. Values with $p<0.05$ and Pearson correlation larger than 0.4 or smaller than -0.4 are deemed significant and shown in bold.
@@
First, let us focus on the results for $d=1$ shown in the top of Figure 3. We observe that degree is not very effective for identifying nodes as for all networks there is only a tiny fraction of unique nodes. However, when moving to count, which additionally accounts for the number of edges in the neighborhood, the uniqueness increases drastically for many networks, sometimes showing results similar to those of $1$-$k$-anonymity. The difference is especially larger in networks with higher degree, less diverse degree distribution, indicated by a higher alpha value, and lower diameter or average distance. Uniqueness obtained by degdist is in many cases equal to $1$-$k$-anonymity, and otherwise very similar. For networks with a high alpha, the difference is larger. This shows that count and degdist appear to be useful for approximating $d$-$k$-anonymity for the networks included. For these, we find that the difference between the two measures is often larger for graphs with a high number of edges, higher median degree, or lower average distance. For the networks with a larger difference, degdist would better approximate $1$-$k$-anonymity.
@@
The vrq measure, which could not fit in our strictness ordering of neighborhood based measures, achieves higher uniqueness than $1$-$k$-anonymity, which very clearly indicates that having knowledge beyond the $1$-neigbhorhood of a target node can have a de-anonymizing effect: it is in many cases even more effective than perfect knowledge of the $1$-neighborhood. We find that the difference is often larger for networks with a high diameter, average distance, or a lower median degree. When looking at hybrid, hence adding knowledge of the $1$-neighborhood structure, we see for only a few networks a slight increase in uniqueness. This implies that the sets of nodes that can be identified with the measures tend to overlap largely, and this additional knowledge of the neighborhood structure only has a small effect on uniqueness.
@@
Turning to the results for $d=2$ at the bottom of Figure 3, we see a very large increase of uniqueness compared to the results for $d=1$. Here the results obtained by degdist are always the same as that of $d$-$k$-anonymity. The count measure achieves similar results to both. vrq always achieves a uniqueness larger than or equal to that of $d$-$k$-anonymity. The generally small difference between vrq and degdist is consistent with the finding in [12] that looking further than the $2$-neighborhood does not have a large effect on uniqueness. Moving to hybrid, we see that this never results in a higher uniqueness when considering the $2$-neighborhood.
@@
Overall, these results demonstrate that having imprecise information, even as little as the number of connections of nodes at distances 1 up to $d$, can be sufficient to uniquely identify a large fraction of nodes. Moreover, this knowledge is more revealing than accounting for complete knowledge using a smaller radius. One may not even need complete structural information to identify many nodes, as it appears that having complete structural information often has a very limited additional effect on uniqueness.
Figure 4: Runtime in seconds (vertical axis) on different empirical network datasets (horizontal axis) when accounting for different measures for $k$-anonymity: degree (lightblue circle), count (blue square), degdist (dark blue circle), $d$-$k$-anonymity (black triangle), vrq (red circle) and hybrid (pink square).

### v2

@@ 4.3 Hybrid
The previously mentioned measures can also be combined, which is done in [77], where the aim is to protect against the so-called $1^{*}$-neighborhood attack. For the corresponding measure, nodes are equivalent if they have isomorphic 1-neighborhoods and if the degree distribution over the neighboring nodes is equal. Hence, this combines the measures of 1-neighborhood isomorphism [11, 74, 73, 12, 66], and VRQ [76, 32]. A parameterized version would combine the notion of $d$-$k$-anonymity with VRQ$(d)$.
4.4 Automorphism based
Several works focus on a very strict scenario in which an attacker should not be able to distinguish between two nodes, even with complete structural information [33, 84, 85]. There are two variants. In [33, 84], the aim is that for each node there should be at least $k-1$ nodes in the same orbit. This implies that the nodes have the exact same structural properties and hence can not be distinguished even when an attacker has perfect structural information of the entire network. In [85], the network is $k$-anonymous if it can be partitioned into $k$ subgraphs that are isomorphic to each other. To achieve $k$-anonymity for these measures, symmetry should be introduced to the network. In particular, for a network to be $k$-anonymous, all its components should be symmetric in at least $k-1$ points, which is not realistic for most real-world network datasets. Anonymizing for this measure would therefore have major impact on the utility of an anonymized version of the network. As this approach is computationally expensive and approximated by the distance-parameterized structural neighborhood measure $d$-$k$-anonymity mentioned in Section 4.2, we do not include automorphism based measures in our comparison.
4.5 Anonymity-cascade
@@
Recent work introduces anonymity-cascade [12]. This algorithm models the scenario where an attacker reuses the nodes that are uniquely identified to identify more nodes in the network. The cascading algorithm starts with all nodes that have a unique $1$-neighborhood structure. Then, it continues in a cascading fashion by identifying neighboring nodes with a unique structure. This way, the reach of the measure is extended beyond the direct neighborhood. This is similar to the attacker scenarios that include a propagation step [86, 87].
4.6 Other measures
@@
For completeness sake, we end with approaches introduced in literature that take into account different structural information than the aforementioned measures. We summarize four such methods.
@@
First, besides vrq as discussed in Section 4.1, the work of Hay et al. [76, 32] introduces two different attacker scenarios for which the first scenario is based on edge facts. This approach measures anonymity by counting the number of candidates given that an attacker knows a number of edges surrounding the node. However, computing all possible subgraphs given a number of edges is very computationally expensive and would not scale to large graphs. At the same time, if the edges are limited to the $d$-neighborhood of the graph, it can be at most as strict as $d$-$k$-anonymity mentioned at the end of Section 4.2.
@@
Second, the hub fingerprints approach assumes that the attacker knows the distance for each node to so-called hub nodes. These are nodes that have a central position in the network and often have a high degree. However, experiments showed that this knowledge helps to identify only few nodes. As this is not an effective attacker scenario, we chose to not include this measure in our comparison.
@@
Third, [88] introduces an attacker scenario where the attacker has additional structural information besides the neighborhood of a node. This work assumes the attacker can obtain information about the centrality of nodes and explicitly aims to protect against this scenario.
@@
Fourth and last, [12] introduces the notion of twin-uniqueness. Under the assumption that the graph is shared, the attacker can identify so-called twin nodes, which are sets of nodes that are connected to the same nodes. When an attacker finds that all candidates for an entity of interest are twins, they can learn all structural information there is to know about the node representing the entity, including the connections. The notion of twin-uniqueness accounts for this phenomenon when computing anonymity and adds the constraint that for a node to be anonymous at least one equivalent node should not be a twin.
@@ 5 Theoretical comparison of $k$-anonymity network measures
In this section, we aim to understand and compare the most prominent measures for anonymity introduced in the literature. We do so by summarizing the measures based on the the structural information they take into account, which allows us to order them based on strictness. We formally define this notion in Definition 1.
Definition 1 (Strictness).
Given two measures $M_{1}$ and $M_{2}$, we say that $M_{1}$ is more strict than $M_{2}$, denoted $M_{1}\geq M_{2}$, if for each pair of nodes $v,w$ it holds that $v\cong_{M_{1}}w$ implies that $\cong_{M_{2}}w$.
This relation defines a partial order on the set of equivalence measures for nodes on a network and has three important implications. First, the more strict measure protects against all attacker scenarios represented by the more lenient measures. Second, given two measures such that $M_{1}\geq M_{2}$, for any graph $G=(V,E)$ it holds that $U_{M_{1}}(G)\geq U_{M_{2}}(G)$: the more strict measure always results in a higher or equal uniqueness, resulting in a lower anonymity. Third, when anonymizing the network for a more strict measure, the minimum number of edge alterations, i.e., node or edge deletion and addition, required for anonymization is at least equal to the number of alterations required when anonymizing for the more lenient measure. We formally prove this in Theorem 1 in Appendix A.
@@
In our comparison, we account for the measures included in Table 3, which shows for each measure its abbreviation and the equation to compute the value for a given node $v$. All measures included, except for degree, are parameterized by parameter $d$. This value determines up to what distance from the considered node the structural information is taken into account. For these parameterized variants, the equivalence classes can be computed recursively [73, 76] by splitting each into new classes as the distance increases. Below, we briefly summarize each of the measures used.
@@
  • count($d$): the number of nodes and edges in the $d$-neighborhood of the considered node.
  • degdist($d$): the degree distribution of the $d$-neighborhood of the considered node. Note that this is a multiset.
@@
  • vrq($d$): the degree of all nodes at distance at most $d$ from the considered node. Note that this is a multiset.
  • hybrid($d$): set containing the results of $d$-$k$-anonymity and vrq($d$), i.e., the exact structure of the $d$-neighborhood of the considered node summarized by its canonical labeling and the degree of all nodes at distance at most $d$ from the considered node.
We can use strictness to order the measures for $k$-anonymity by comparing them based on their reach, and what structural information they take into account. First, we look at the neighborhood based measures being degree, count, degdist and $d$-$k$-anonymity. As illustrated in the second and fourth column of Table 2, these measures all have the same reach, namely the $d$-neighborhood of the node. However, the measures do differ in the extent to which they capture particular structural properties. For vrq, the reach is larger, but the structural information is less precise. As a result, vrq can not be included in the strictness ordering of neighborhood based measures. As the hybrid measure combines both $d$-$k$-anonymity and vrq, this measure is more strict than both. We formally prove that this ordering holds by showing that the stricter measures capture properties of the less strict measures. The ordering resulting from Theorems 2 and 3, which can be found in Appendix A, is visualized in Figure 1.
Figure 1: Ordering of anonymity measures based on strictness. An arrow between two measures ($A\rightarrow B$) implies measure $A$ is more strict than measure $B$. For readability, arrows implying strictness through transitivity are not included.
@@ 6 Empirical comparison of k-anonymity measures
In this section, we empirically compare the anonymity measures in Table 3 on a wide range of network datasets. First, we summarize the experimental setup. Second, we show how different measures affect the anonymity by comparing uniqueness. Third, we study the effect of a cascading step on anonymity. Fourth, we investigate how the choice in measure affects the trade-off between anonymity and data utility when anonymizing the network. Lastly, we assess the runtime required to compute anonymity with each measure.
Table 4: Real-world network datasets used in the experiments. For each network, we list the number of nodes, edges, average degree, median degree, largest degree, clustering coefficient, assortativity, diameter, average distance and power law exponent $\alpha$.
Network $|V|$ $|E|$
Avg.
@@
6.1 Experimental setup
For the experiments, we use the networks in Table 4, which can each be found in their corresponding repositories. All networks are interpreted as undirected, and additional properties such as weights or timestamps are ignored. We assess anonymity in terms of uniqueness, i.e., the fraction of nodes that are $1$-anonymous, using the measures listed in Table 3 and a more generic version of anonymity-cascade as discussed in Section 4.5. We report on results for both $d=1$ and $d=2$, as previous work found a relatively large increase in uniqueness for $d$-$k$-anonymity with $d=2$ with only relatively small differences compared to larger values of $d$ [12]. To compute uniqueness, we use the same computational approach as described in [73]. Code used can be found on github [^1] https://github.com/RacheldeJong/ANONET.
To assess the anonymity vs. utility trade-off, we use edge sampling [11] as anonymization algorithm, which removes edges at random. The resulting graphs are created by repeatedly deleting 1% of the edges, resulting in 101 networks for each run including the initial network. For each of these graphs, utility metrics as discussed in Section 3.1 are computed using igraph [96]. Data utility is assessed in terms of performance on three common downstream tasks in network analysis listed below:
  • Robustness: the fraction of nodes remaining in the largest connected component after anonymization.
  • Community structure: the NMI [62] of communities found before and after anonymization. Communities are found by 10 runs of the Leiden algorithm [61] followed by a consensus clustering step [62].
  • Node centrality: the overlap of nodes in the top 100 most central nodes according to betweenness centrality [63] before and after anonymization.
To account for non-determinism of the edge sampling algorithm, the anonymization process and utility computation for all 101 obtained networks are repeated five times. Utility, uniqueness and runtime values are averaged over these five runs with a time limit of 3 hours. Results $\pm$ one standard deviation are reported.
@@ 6.2 Comparing uniqueness
From Section 5, we know that uniqueness, the inverse of anonymity, is expected to be higher for stricter measures and lower for less strict measures. Figure 2 shows how the uniqueness varies in empirical networks by reporting this value for each measure at both $d=1$ and $d=2$. In Appendix B, we report for each measure the difference with the previous (less strict) measure in the legend of Figure 2. In an attempt to understand performance differences, we compare each difference with various topological network properties, of which the results can be found in Table 5. This table shows for each combination of difference and network property the Pearson correlation and corresponding $p$-value. Plots of the values for all significantly correlated combinations can be found in Appendix C. We use these findings in the discussion of the results.
Table 5: Correlation between network properties (leftmost column) and the difference observed between the use of two measures (second up to last column). For each combination of network property and outcome both the correlations and $p$-value are reported. Values with $p<0.05$ and Pearson correlation larger than 0.4 or smaller than -0.4 are shown in bold.
@@
First, we focus on the results for $d=1$ shown in the top of Figure 2. We observe that degree is not very effective for identifying nodes as for all networks there is only a tiny fraction of unique nodes. However, with count, which additionally accounts for the number of edges in the neighborhood, the uniqueness increases drastically for many networks, sometimes showing results similar to those of $1$-$k$-anonymity. The difference appears to be especially larger in networks with higher degree, lower diameter or average distance, and a more balanced degree distribution, indicated by a higher alpha value (see Table 4). Uniqueness obtained by degdist is in many cases equal to $1$-$k$-anonymity, and otherwise very similar. For networks with a high alpha, the difference is larger. This shows that count and degdist, measures that use imprecise structural information and hence represent a weaker attacker scenario, appear to be useful for approximating $d$-$k$-anonymity, which accounts for exact structural information.
@@
The vrq measure, which could not fit in our theoretical strictness ordering of neighborhood based measures, achieves higher uniqueness than $1$-$k$-anonymity, which indicates that having knowledge beyond the $1$-neighborhood of a considered node can have a strong de-anonymizing effect: it is in many cases even more effective than perfect knowledge of the $1$-neighborhood. We find that the difference is often larger for networks with a high diameter or average distance, or lower median degree. For the hybrid measure, adding exact knowledge of the $1$-neighborhood structure to vrq, we see for only a few networks a slight increase in uniqueness. This implies that the sets of nodes that can be identified with the measures tend to overlap largely, and this additional knowledge of the neighborhood structure only has a small effect on uniqueness. Hence, when choosing an attacker scenario to protect against, vrq may be more effective than $d$-$k$-anonymity.
@@
Results for $d=2$ at the bottom of Figure 2 show a very large increase of uniqueness compared to the results for $d=1$. Here the uniqueness values obtained by degdist are always the same as that of $d$-$k$-anonymity and that of count is similar to both. vrq always achieves a uniqueness larger than or equal to that of $d$-$k$-anonymity. The generally small difference between vrq and $d$-$k$-anonymity with $d=2$ is consistent with the finding in [12] that accounting for information beyond the $2$-neighborhood does not have a large effect on uniqueness. For the hybrid measure, we see that this never results in a higher uniqueness when considering the $2$-neighborhood.
@@
Overall, these results demonstrate that having imprecise information, even as little as the number of connections of nodes at distances 1 up to $d$, can be sufficient to uniquely identify a large fraction of nodes. Moreover, this knowledge is more revealing than complete knowledge using a smaller radius. One may not even need complete structural information to identify many nodes, as it appears that having complete structural information often has a very limited additional effect on uniqueness.
6.3 Anonymity-cascade
The results from the previous section showed that measures that reach further lead to a much higher uniqueness and hence represents more effective attacker scenarios. In this section, we investigate the effect on uniqueness when using a different method to increase, namely by cascading the information of unique nodes to identify more nodes. To do so, we measure uniqueness using a generic variant of anonymity-cascade [12], explained in Section 4.5. The algorithm consists of two steps. First, we identify nodes using the initial measure. Second, we use the cascade measure to identify more nodes in a cascading fashion. This is also illustrated in Figure 3. The cascading process can continue for multiple levels. If only one level of cascading is used, we refer to this as $C_{1}$. The process can then repeat until no more nodes can be identified, referred to as cascading final $C_{f}$.
Figure 4: Uniqueness using anonymity-cascade one level ($C_{1}$). Each figure corresponds to a different initial measure (triangle and line) and is combined with all of the cascade measures (dots). The grey bar indicates the highest obtained uniqueness for the network given the initial measure. Measures used are: degree (lightblue), count (blue), degdist (dark blue), $d$-$k$-anonymity (black triangle), vrq (red) and hybrid (pink).
The combinations of initial and cascade measure consist of all 36 combinations of the six measures in Table 3. This allows us to investigate all combinations of attacker knowledge. Besides the cases where the levels of knowledge are equal, this also includes the cases where an attacker has little starting knowledge and more local knowledge surrounding the identified nodes, hence more cascade knowledge. While we fully acknowledge that not every combination of measures is equally realistic, we chose to include all to observe the cascading effect of further reach on the uniqueness. We focus on one level of cascading, as this is a more realistic scenario. (The results for cascading final can be found in Appendix D)
Figure 4 shows the uniqueness for anonymity-cascade with one level of cascading. Each subfigure shows results for a different initial measure, denoted by the line with a triangle, and results of one level of cascading, denoted by the colored dots, using each of the six measures as cascading measure. The gray bar is added for readability of the results and shows the highest uniqueness obtained.
First, the results for degree show a very low initial uniqueness. However, one level of cascading results in a much higher uniqueness, indicating that even with minimal initial knowledge, i.e., the degree of a node, many more nodes can be identified with additional cascade knowledge. When using degree as cascading measure, this impact is often smaller. In most cases using vrq as cascading measure has the largest impact on the uniqueness.
The observed uniqueness when using degree as cascading measure, still differs substantially from results for the other measures. Results show that count, degdist and $d$-$k$-anonymity achieve similar results, where using vrq as cascade measure does increase uniqueness. With vrq as initial measure we do observe both a higher initial uniqueness and uniqueness after cascading.
Overall, these results show, similar to the results of Section 6.2, that by reaching further, with one or more (see Appendix D) cascading steps, the uniqueness increases immensely. This effect is especially clear when using degree as initial measure: if the nodes with unique degrees can be identified, this can be reused to identify many other nodes.
Figure 5: Pareto optimal solutions found in terms of uniqueness (horizontal axis) and performance on a downstream task (vertical axis). Each dot represents a solution found (partially anonymized network) when anonymizing using edge sampling for one of the five anonymity measures (colors). Error bars indicate results $\pm$ one standard deviation.
6.4 Data utility
In this section, we empirically investigate how the chosen measure affects the trade-off between utility and anonymity when anonymizing the networks by means of edge sampling [11]. Due to computational reasons, we perform anonymization on the 14 smallest networks. For each partially anonymous network found during the anonymization process, we determine the uniqueness, difference in robustness, community structure, and 100 most central nodes. To represent the trade-offs found when using different measures, Figure 5 contains for two networks the Pareto fronts which represent the best trade-off values in terms of observed uniqueness and difference in performance on the three downstream tasks. Hence, each result (dot) represents a network with a different percentage of deleted edges. Figure 5 shows the results on two networks. Results for the remainder of the networks, which appear to follow similar behavior, can be found in Appendix E.
Overall, we can see a large difference between the obtained Pareto fronts for the measures where the most lenient measures, degree and count, find better trade-offs and the measure with the largest starting uniqueness, vrq, finds worse trade-offs. When using the degdist and $d$-$k$-anonymity measures, the solutions found are very similar in terms of uniqueness and utility. Surprisingly, for the count measure, the trade-offs found are clearly better. This difference is larger for networks where the starting uniqueness differs more between the count and degdist measure, such as “FB Reed98”.
Comparing the downstream tasks, we observe that for both community structure and node centrality a big decrease in utility needs to be sacrificed to improve uniqueness. For the largest component size, this difference is more gradual, yet steep for vrq. Hence, these results show that when using a more strict measure, or one that reaches further, more utility has to be sacrificed in order to obtain the same uniqueness.